# GlucoLens: Explainable Postprandial Blood Glucose Prediction from Diet and Physical Activity

### v1

@@ II-E Hyperglycemia Detection and Counterfactual Explanations
Models RF, Ridge, MLP
@@
Random seeds 3 different seeds
Total # experiments $3\times 5\times(3+3+13)\times 3=855$
TABLE III: Variations of the multilayer perceptron (MLP regressor). Different variations have been tested by varying the depth and size of each layer.
Variation # Hidden Sizes of
@@
III Results
Figure 4: Overall Pearson correlation matrix with the coefficients multiplied by 10 to avoid printing decimal points in the figure. The amount of fiber, duration of time in stepping, and total fat intake are some of the features working in favor of lowering postprandial area under the curve (PAUC) and related outcomes. PAUC is represented as ‘absolute_auc’ in this figure and increase in PAUC after lunch, $\Delta$PAUC as ‘respective_auc’. On the other hand, duration spent sitting, higher fasting glucose, a higher BMI, and working at home are some of the features positively correlated with PAUC.
III-A Best Combination for Each Outcome
Three different outcomes were chosen for the regression problem: i) PAUC, ii) $\Delta$PAUC, and iii) MaxPBG. In Table IV, we see that for PAUC, $\Delta$PAUC, and MaxPBG, the best NRMSE values are 0.123, 0.622, and 0.132 respectively. RF predictors obtained all three best results. Although 13 different variations were tested for MLPs, none of them achieved the best performance on any of the three tasks.
TABLE IV: Summary of the best results for each target outcome. Three different models were tested with five different combinations. Meanings of the outcomes PAUC, $\Delta$PAUC, and MaxPBG can be found in Fig. 1. RF = random forest. NRMSE = normalized root mean squared error.
Target Best Best Best Test Test
outcome feature set model hyperparameters RMSE NRMSE
PAUC All RF $n_{est}=10$ 39.2 0.123
$\Delta$PAUC Self + Macro RF $n_{est}=100$ 58.8 0.622
MaxPBG All RF $n_{est}=100$ 16.4 0.132
III-B Effect of Feature Sets
The impact of different input features and their combinations is described in Table V. All features and Sensor + Macro result in the best NRMSE scores ($0.123$ and $0.132$) across all feature sets for estimating PAUC and MaxPBG, while the best NRMSE ($0.622$) for $\Delta$PAUC estimation comes from the Self + Macro feature set. The average NRMSE across all five feature sets are $0.130$, $0.736$, and $0.144$ for PAUC, $\Delta$PAUC, and MaxPBG estimation, respectively.
TABLE V: Effect of feature set on the regression tasks. The corresponding hyperparameters of the random forest (RF) and Ridge regression models are presented in parentheses. NRMSE = normalized root mean squared error.
Feature PAUC $\Delta$PAUC MaxPBG Model Model Model
set NRMSE NRMSE NRMSE for PAUC for $\Delta$PAUC MaxPBG
Sensor + Macro 0.123 0.798 0.132 RF (50) RF (100) RF (100)
Self + Macro 0.139 0.622 0.161 RF (10) RF (100) Ridge (0.01)
Sensor + GL 0.125 0.827 0.134 RF (100) RF (100) RF (100)
Self + GL 0.139 0.626 0.159 Ridge (0.01) RF (50) Ridge (0.01)
All 0.123 0.806 0.132 RF (10) RF (100) RF (100)
Average 0.130 0.736 0.144 - - -
III-C Performance of MLP
Out of the 13 different variations of MLP regressors, none of them were the best for any of the three outcomes or any of the 15 outcome-feature_set combinations, as shown in Table IV and Table V. It encouraged us to look deeper into the results of the MLP variations. We present the best performances by each of the 13 MLP variants for each of the three outcomes in Table VI. MLP variation no. 13 is the largest of all the MLP variations used in our experiments and it performed better than any other MLPs in all three tasks. It raises the question of whether even larger and deeper models would be more accurate and potentially better than the RF and Ridge models. We wish to investigate this in future work.
TABLE VI: Percentage of error in terms of NRMSE (i.e. NRMSE $\times$ 100) for different variations of MLP in three different regression tasks. Results are rounded to the nearest integer unless any other result rounds to the same integer as the best result. The details of the MLP variations can be found in Table LABEL:tbl:mlp_hyperparams. For the best results of PAUC, $\Delta$PAUC, and MaxBPG, the corresponding feature sets are Sensor + GL, Self + Macro, and Sensor + GL.
Variation 1 2 3 4 5 6 7 8 9 10 11 12 13
PAUC 74 29 27 22 21 19 18 18 19 18 18 18 17
$\Delta$PAUC 89 82 81 81 80 80 81 79 80 79 79 80 78
MaxPBG 50 24 24 21 21 19 18.3 18.0 19 18 19 19 17.5
III-D Prediction with XGBoost and TabNet
One may wonder if there are other machine learning models or neural networks that could be used to evaluate the system’s efficacy. We believe that advanced neural networks integrated with our system will be able to perform more accurate results but it will require more data to train them. To explore more, apart from the random forest, MLP, and ridge regressions, we also trained two comparatively more advanced ML models: XGBoost[25] and Tabnet[26]. In our experiments, Random Forest outperformed both TabNet and XGBoost in estimating PAUC. TabNet trained with 100 epochs provided an NRMSE of 0.147 and XGBoost provided an NRMSE of 0.137. A summary of the performance of XGBoost and TabNet for different feature sets is presented in Table VII.
@@ III-D Prediction with XGBoost and TabNet
TABLE VII: Test set normalized root mean squared errors (NRMSE) of the XGBoost and TabNet models for different feature sets for estimating the postprandial area under the curve (PAUC).
Feature set TabNet XGBoost
Sensor + GL 0.160 0.137
Sensor + Macro 0.147 0.139
Self + GL 0.154 0.152
Self + Macro 0.151 0.149
All 0.151 0.137
III-E Correlation Analysis
In addition to the results of the prediction problems, we also present Pearson correlations of the factors with the outcome variable PUAC in Fig. 4. As it is recommended to keep MaxPBG below 140 mg/dL for non-diabetic people [27], a binary outcome of MaxPBG $>$ 140 was added to the correlation analysis to find out what features are positively or negatively correlated with this outcome. The observed correlations of this outcome with the input features were similar to the correlations of the other three outcomes.
We notice a positive correlation between the negative outcomes (high PAUC or high MaxPBG) and the features: BMI, sitting duration, and working at home. We also notice that overall, a delayed lunch is correlated with a lower AUC. Finally, fiber is negatively correlated with the outcomes as expected. This correlation analysis also shows us the risk of postprandial hyperglycemia because of higher PAUC or $\Delta$PAUC metrics. If we look at Fig. 4, especially the row of ‘MaxPBG $>$ 140’, which directly represents the risk of hyperglycemia, we notice that it has a very high correlation coefficient with both ‘abosolute_auc’ (PAUC) and ‘respective_auc’ ($\Delta$PAUC), with values 8 and 5 respectively (on a scale of -10 to 10). It indicates that a higher PAUC or $\Delta$PAUC may lead a person to develop hyperglycemia in the long run.
(a) Original label: Hyperglycemia, CF label: Normal
(b) Original label: Normal, CF label: Hyperglycemia
Figure 5: Counterfactual explanations for a normal and a postprandial hyperglycemic event. Features not presented in the figures had the same value in the original and all its counterfactual examples. Feature values of the counterfactual examples were plotted with small shifts (+0.01 for CF1, +0.02 for CF2, +0.03 for CF3) for better visualization. Units not shown in the figure: stepping_total (s), sitting_at_work (s), fasting_glucose (mg/dL), recent_cgm (mg/dL), lunch_time (h), work_start_time (h). Plotted values are scaled. Multiply by the given factors to obtain actual values.
III-F Results of Hyperglycemia Detection and Diverse Counterfactual Explanations
GlucoLens achieves an accuracy of 74% and an F1 score of 0.55 in hyperglycemia detection with a small training dataset. The visualization of counterfactual explanations in Fig. 5 highlights insights for avoiding hyperglycemia by examining the influence of various features on postprandial blood glucose outcomes. The original data point of Fig. 5(a) associated with hyperglycemia (Outcome=1) is represented by a solid blue line with a blue dot, while three counterfactual scenarios are depicted using dashed lines. Counterfactual 1 (orange, Outcome=0) illustrates a scenario where increasing fiber intake changes the outcome. In Counterfactual 2 (green, Outcome=0), increasing fiber by a large amount, despite additional cholesterol intake, changes the class label. Counterfactual 3 (red, Outcome=0) does not make any change in diet but increases physical activity (stepping duration) significantly to change the outcome.
Similarly, in the second example with a normal outcome (Outcome=0) as shown in Fig. 5(b), counterfactual scenarios reveal potential pathways to hyperglycemia. Counterfactual 1 (orange, Outcome=1) highlights how an increase in sedentary time at work (in seconds) and an early work start time can increase the chance of hyperglycemia. In Counterfactual 2 (green, Outcome=1) and Counterfactual 3 (red, Outcome=1) a higher starting fasting glucose and recent CGM readings in addition to much longer sedentary duration contributed to the class label to flip from normal to hyperglycemia. These findings highlight the importance of diet, activity, and habits in managing glucose levels, demonstrating the value of counterfactual explanations for personalized recommendations.
IV Discussion
IV-A Summary of the Results
@@
We have presented an important problem of estimating the postprandial area under the curve (PAUC), change in PAUC ($\Delta$PAUC), and maximum postprandial blood glucose level (MaxPBG). Based on our experiments, we found that random forest (RF) models outperform multilayer perceptions (MLP) and ridge regression models as shown in Table V. We also chose our features from five different feature sets as illustrated in Fig. 3. Although our feature sets use 31 features as input to the model when choosing the set of ‘All’ features, all these features are derived from only four sources of data, which makes it easier to have the required data easily available when needed for the model.
IV-B Different Interpretation of the Results
@@
Our best-found NRMSE of 0.123 implies that the predicted PAUC value was on average within a 12.3% error margin from the actual values of PAUC. To interpret the performance in another way, we also explored the percentage of test cases falling within an error tolerance. From the actual PAUC values and the predicted PAUC values, the ratio of test cases within 5%, 10%, 15%, and 20% errors were calculated. Among the predictions made by our system, we have verified that 33% of the cases had an error of less than 5%, 61% of the cases had an error of less than 10%, 81% of cases had an error of less than 15%, and 93% cases had an error less than 20%. Thus only 19% of the cases had an error more than 15% and only 7% cases had an error over 20%. We believe that with more training data, our system will be able to predict most of the cases within a very small margin of error. The percentage of test cases within 5%, 10%, 15%, and 20% error margins for the XGBoost model were 30%, 55%, 78%, and 92% respectively, all lower than our random forest model. Therefore, our random forest model outperforms XGBoost in this metric as well.
IV-C Alternative Feature Search Methods

### v2

@@ 2.4 ML Models
RF, Ridge, MLP, XGBoost, TabNet, GPT-3.5, GPT-4, Mistral Large,
Predictors Gemini Flash 2.0, Claude Opus 4, Grok 3, Deepseek V3, Gly_Hybrid,
Gly_Hybrid_v2, Gly_Max, Hybrid Predictors for Classification
(RF+MLP, RF+XGB, XGB+MLP, RF+XGB+MLP).
@@
Table 3: Variations of the multilayer perceptron (MLP regressor). Different variations have been tested by varying the depth and size of each layer.
Variation # Hidden # Nodes in
@@
Figure 5: A sample zero-shot prompt for the LLMs. The input field is varied for each different example; everything else in the prompt remains the same.
Five different architectures: Random Forest (RF), Ridge Regressor, MLP Regressor, XGBoost, and TabNet, were trained as potential backbone architectures of the GlucoLens system. MLPs and Random Forests are capable of making effective predictions with proper feature extractions Azghan et al. (2025); Mamun et al. (2023), whereas linear regression methods with regularization, such as Ridge Regression McDonald (2009) and Lasso Regression Ranstam and Cook (2018), are popular choices as comparison models. Neural networks are usually data-hungry, which is why, on small datasets, classical ML methods can show competitive and even better performance than neural networks.
Additionally, zero-shot large language models were used for LLM-only predictions based on the feature values. The structure of the prompt used for the LLMs is presented in Fig. 5. The content for the "Input:" field is changed for each different datapoint, whereas the rest of the body of the prompt remains the same. In the next step, hybrid predictors that use both zero-shot LLM predictions and a trainable regressor (RF or XGBoost) were implemented. Three variations of the hybrid predictors were implemented and trained. The first variation (Gly_Hybrid) uses the existing features and adds the predictions by six advanced LLMs as additional inputs. The second hybrid variation (Gly_Hybrid_v2) uses the existing features and the prediction by only the best-performing LLM (in our case, Claude Opus 4) for training and testing. Finally, the third hybrid variation (Gly_Max) uses a Gaussian noise-based augmented feature set based on existing features, predictions by Claude Opus 4, and a trainable regressor.
Three variations of RF, three variations of Ridge, and thirteen variations of MLP, one version of XGBoost, and one version of TabNet were trained and tested for the Gly_Base version of the GlucoLens system. The different models and variations for AUC, MaxBGL, and Hyperglycemia prediction are presented in Table 2 and Table LABEL:tbl:mlp_hyperparams. For AUC and MaxBGL predictions, 20% data instances were assigned to the test set, and 80% data instances were assigned to the training set. For hyperglycemia detection, six different variations of train-test split were used, which are discussed in details in Section 3.4 and Table 7.
2.5 Hyperglycemia Detection and Counterfactual Explanations
Figure 6: The pipeline of classification of hyperglycemia with GlucoLens system’s hybrid classifier. A soft voting based on the probabilities of classes suggested by the RF, XGBoost, and MLP (version 13 of Table LABEL:tbl:mlp_hyperparams) is used to make the final predictions. This hybrid method outperforms the prediction performances of single classifiers. The test set contains only real datapoints, so that the evaluation is not biased by the data generation method, whereas the balanced training data contains both real and synthetic datapoints.
GlucoLens was also trained for hyperglycemia detection with RF, XGB, MLP, backbone models, as well as hybrid combinations of these three, as shown in Fig. 6. More details about all the different variations of the models and different splits of training and test sets are described in Section 3.4. Counterfactual explanations can provide insights on features responsible for an undesired health outcome and possible remedies to overcome them Mamun et al. (2024); Arefeen et al. . As a primary goal of the paper was to explore more knowledge about the reasons for and ways to prevent hyperglycemia, a DiCE-based counterfactual explanations generator is integrated with GlucoLens. GlucoLens provides multiple counterfactual explanations that are diverse and achievable with small perturbations from the original example Mothilal et al. (2020).
3 Results
3.1 Prediction of AUC with Different Backbone Architectures
The first problem we want to solve is the prediction of postprandial AUC and we used five different architectures as the backbone of our model, that includes 3 different versions of Random Forest (RF), 3 different versions of Ridge regressors, 13 versions of MLP, and finally, XGBoost Chen and Guestrin (2016) and TabNet Arik and Pfister (2021). According to our experiments, GlucoLens performs the best with an RF backbone with a normalized root mean squared error (NRMSE) of 0.123. TabNet trained with 100 epochs provided an NRMSE of 0.147, and XGBoost provided an NRMSE of 0.137. A summary of the performance of all these different backbones can be seen in Table 4. We believe that advanced neural networks integrated with our system will be able to produce more accurate results, but it will require more data to train them. We show in Section 3.4, where the results of hyperglycemia detection are discussed, that the performance on the test set monotonically improves as we increase the size of the training dataset.
@@ 3.1 Prediction of AUC with Different Backbone Architectures
Table 4: Normalized Root Mean Squared Errors (NRMSE) of our GlucoLens models (RF, Ridge, MLP, XGBoost, TabNet) for different feature sets in the prediction of postprandial AUC. Explanations of the feature sets can be found in Table 1.
Feature Set RF Ridge MLP XGBoost TabNet
Sensor + GL 0.125 0.139 0.169 0.137 0.160
Sensor + Macro 0.123 0.142 0.172 0.139 0.147
Self + GL 0.142 0.139 0.178 0.152 0.154
Self + Macro 0.139 0.142 0.172 0.149 0.151
All 0.123 0.140 0.176 0.137 0.151
3.2 Prediction with LLMs
The results presented in Table 5 and Fig. 7 showcase the performance of various configurations of our proposed solution. These configurations include Gly_Base, which does not utilize large language models (LLMs); Gly_LLM, a zero-shot prediction approach using only LLMs; Gly_Hybrid, combining LLM-based predictions with our system; Gly_Hybrid_v2, which integrates predictions exclusively from the best-performing LLM (Claude); and Gly_Max, an enhanced version of Gly_Hybrid_v2 with augmented training data incorporating Gaussian noise. Performance of the hybrid models was evaluated by integrating LLMs with one of the two best-performing regressor backbones: Random Forest (RF) or XGBoost. Among the configurations, Gly_Base with RF backbone achieved the best performance. Interesting findings can be observed in Fig. 7, where we can see that Claude Opus 4 outperformed all other LLM-based models by a large margin. These findings indicate that it may not be wise to ask LLMs for personalized meal plans as they cannot accurately predict the consequences of a certain combination of diet and other factors on the postprandial blood glucose levels. We used the ‘All’ feature set for the LLM-based experiments.
Table 5: AUC NRMSE results of different variations of our solution. Gly_Base = GlucoLens regressor with no LLM, Gly_LLM = LLM only prediction (zero-shot) after multimodal data processing by GlucoLens. The hybrid predictors are explained in Table 2.
Backbone Gly_Base Gly_LLM Gly_Hybrid Gly_Hybrid_v2 Gly_Max
RF 0.123 0.290 0.241 0.238 0.226
XGBoost 0.137 0.236 0.242 0.259
Figure 7: Normalized root mean squared errors on the prediction of AUC with different LLMs and hybrid classifiers of our system. Explanations of Gly_Base and the hybrid predictors can be found in Table 2. Explanations of the feature sets can be found in Table 1.
3.3 Performance of MLP and Prediction of MaxBGL
Two different outcomes were chosen for the regression problem of predicting AUC and MaxBGL. In Fig. 8, we see that for AUC and MaxBGL, the best NRMSE values are 0.123 and and 0.132, respectively. RF predictors obtained all three best results. Although 13 different variations were tested for MLPs, none of them achieved the best performance on any of the three tasks.
Figure 8: Summary of the best results for each target outcome with RF, Ridge, and MLP regressors. Three different models were tested with five different combinations. Meanings of the outcomes AUC and MaxBGL can be found in Fig. 1. RF = random forest. NRMSE = normalized root mean squared error.
Figure 9: Normalized root mean squared errors (NRMSE) for different variations of MLP in two different regression tasks. Best NRMSE for AUC prediction is 0.169, and best NRMSE for MaxBGL prediction is 0.175. The details of the MLP variations can be found in Table LABEL:tbl:mlp_hyperparams. The best performance among MLPs was achieved with the MLP variation no. 13 and the feature set ‘Sensor + GL’.
Out of the 13 different variations of MLP regressors, all of their NRMSE metrics were higher than the corresponding RF regressors, as shown in Fig. 8. It encouraged us to look deeper into the results of the MLP variations. We present the best performances by each of the 13 MLP variants for both outcomes in Fig. 9. MLP variation no. 13 is the largest of all the MLP variations used in our experiments, and it performed better than any other MLP versions in both tasks. It raises the question of whether even larger and deeper models would be more accurate and potentially better than the RF and Ridge models. We wish to investigate this in future work.
Figure 10: Counterfactual explanations for a normal and a postprandial hyperglycemic event. Features not presented in the figures had the same value in the original and all its counterfactual examples.
3.4 Results of Hyperglycemia Detection and Diverse Counterfactual Explanations
Table 6: GlucoLens system’s classification results with different pure and hybrid backbones for hyperglycemia detection on the 87% training and 13% test split. All metrics are averages over 100 repetitions with different random seeds.
Classifier Accuracy Precision Recall F1
RF 0.698 0.737 0.699 0.685
XGB 0.685 0.720 0.692 0.682
MLP 0.620 0.626 0.620 0.589
RF+XGB 0.695 0.730 0.695 0.683
RF+MLP 0.668 0.700 0.668 0.650
XGB+MLP 0.687 0.712 0.687 0.672
RF+XGB+MLP 0.712 0.740 0.712 0.702
Table 7: Results of the GlucoLens hyperglycemia detection system with RF+XGB+MLP based hybrid backbone as we increase the training data size. All metrics are averages over 100 repetitions with different random seeds. An improvement of the performance metrics can be observed except for the last row, when only 1% of the dataset is withheld for testing. In that case, the evaluation in any trial is vulnerable to producing 0 values for precision, recall, or F1 score, as there is only 1 example from each class in the test set. Eventually, it affects the overall average of those metrics.
Size of training set Accuracy Precision Recall F1
70% training, 30% test 0.674 0.706 0.674 0.660
80% training, 20% test 0.660 0.729 0.702 0.690
87% training, 13% test 0.712 0.740 0.712 0.702
90% training, 10% test 0.717 0.744 0.717 0.705
95% training, 5% test 0.733 0.751 0.733 0.716
99% training, 1% test 0.730 0.625 0.730 0.660
GlucoLens achieves an accuracy of 73.3% and an F1 score of 0.716 in hyperglycemia detection with a small dataset of 159 data points, and we prove with experiments that the performance improves as we increase the size of the dataset, as seen in Table 7. RF, XGB, MLP, and their combinations were first trained and evaluated for a split of approximately 87% training and 13% test set. This split was created by randomly separating 10 positive class (hyperglycemia) examples and 10 negative class (normal) examples to create the test set. The rest of the examples were used to create the training set. Thus, the test set was already balanced, and the training set was balanced with ADASYN He et al. (2008). The results for different classifiers can be seen in Table 6. We notice that the hybrid classifier that uses soft voting among an ensemble of RF, XGB, and MLP outperform the other combinations of classifiers. Note that the performance improvement in ensemble classifiers over individual classifiers are not due to the increased number of parameters because each model of the ensemble classifiers is trained independently (on the same training dataset), and only during inference on the test dataset, they work together to make more reliable predictions. We used the ‘All’ feature set for the hyperglycemia detection experiments.
In the next part of the experiments, we show that as we increase the size of the training set, the performance keeps increasing. We varied the size of the training set from 70% of the total dataset to up to 99% of the dataset, and we noticed a strictly increasing pattern in all four metrics until the training set is 95% of the dataset, as seen in Table 7. This supports the fact that with a bigger dataset, the performance of our method can be much higher than what we are getting with this small dataset. For the last case, when 99% of the dataset is used for training, the test set consists of only 1 example from each class. This case, makes the evaluation of the test set highly vulnerable for the precision and recall metrics, as unless the classifier predicts the whole test set perfectly in any trial, one of these two metrics would be zero for that trial, which would affect the F1 score as well. However, we believe that in a significantly larger dataset, even 1% of the dataset could be used for testing. Overall, the best performance was found with RF+XGB+MLP with 95%, 5% train-test split with an accuracy of 73.3% and a macro average F1 score of 0.716 with a precision of 0.751 and a recall of 0.733.
The visualization of counterfactual explanations in Fig. 10 highlights insights for avoiding hyperglycemia by examining the influence of various features on postprandial blood glucose outcomes. The first example of Fig. 10 represents a hyperglycemic case, characterized by a fiber intake of 1 gram and a stepping duration of 8.95 minutes. Counterfactual scenarios suggest pathways to achieve normal blood glucose levels. Option 1 illustrates that increasing fiber intake to 5 grams can lead to a normal outcome. Option 2 shows that substantially increasing stepping duration to 39.38 minutes also results in a normal blood glucose level. These scenarios emphasize the impact of dietary and physical activity modifications in managing glucose levels.
In contrast, a second example in Fig. 10 represents a normal blood glucose level with current feature values of a work start time of 11 PM, sitting duration of 48.31 minutes, lunch time at 1 PM, and a lunch calorie intake of 780 kCal. Counterfactual scenarios reveal potential transitions to hyperglycemia. The counterfactual example generated against that example demonstrates how starting work earlier at 6 AM, increasing sitting duration to 148.62 minutes, shifting lunch to 12 PM, and increasing lunch calories to 827 kCal can contribute to hyperglycemia. These findings highlight the importance of balancing work habits, meal timing, and physical activity for glucose management, underscoring the value of counterfactual explanations for personalized recommendations.
4 Discussion
4.1 Summary of the Results
@@
We have presented an important problem of estimating the postprandial area under the curve (AUC), the maximum postprandial blood glucose level (MaxBGL), and detecting hyperglycemia. Based on our experiments, we found that random forest (RF) models outperform multilayer perceptrons (MLP), ridge regression, XGBoost, and TabNet models in AUC prediction as shown in Table 4. We also chose our features from five different feature sets as illustrated in Table 1. Although our feature sets use 31 features as input to the model when choosing the set of ‘All’ features, all these features are derived from only four sources of data, which makes it easier to have the required data easily available when needed for the model.
Our LLM-based experiments suggested that Claude Opus 4 outperforms GPT 3.5 Turbo, GPT 4, Gemini 2.0 Flash 001, DeepSeek V3, and Grok 3 in AUC predictions. We also found that hybrid models outperform LLM-only models, whereas augmented training data with Gaussian noise further improves the performance of the hybrid models. However, fully trainable base regressor backbones of our GlucoLens system outperform both LLM-based and hybrid versions of the predictors for AUC prediction.
Finally, in hyperglycemia prediction, we find that GlucoLens with ensemble classifiers with RF, XGB, and MLP outperform the system with individual classifiers, achieving an accuracy of 73.3% and an F1 score of 0.716, as shown in Table 6 and Table 7. Additionally, we empirically prove that the GlucoLens hyperglycemia prediction system monotonically improves its test set performance as we increase the size of the training data.
4.2 Different Interpretation of the Results
@@
Our best-found NRMSE of 0.123 implies that the predicted AUC value was on average within a 12.3% error margin from the actual values of AUC. To interpret the performance in another way, we also explored the percentage of test cases falling within an error tolerance. From the actual AUC values and the predicted AUC values, the ratio of test cases within 5%, 10%, 15%, and 20% errors were calculated. Among the predictions made by our system, we have verified that 33% of the cases had an error of less than 5%, 61% of the cases had an error of less than 10%, 81% of the cases had an error of less than 15%, and 93% cases had an error of less than 20%. Thus only 19% of the cases had an error of more than 15%, and only 7% of the cases had an error over 20%. We believe that with more training data, our system will be able to predict most of the cases within a very small margin of error. The percentages of test cases within 5%, 10%, 15%, and 20% error margins for the XGBoost model were 30%, 55%, 78%, and 92%, respectively, all lower than our random forest model. Therefore, our random forest model outperforms XGBoost in this metric as well.
4.3 Alternative Feature Search Methods